OpenStudy (anonymous):

prove the following trig identity 2/(sqrt 3 cos x +sin x)= sec((pi/6)-x) help please!!!

OpenStudy (anonymous):

......please

OpenStudy (anonymous):

Is it ? \[\frac{2}{\sqrt{3} \cos(x) + \sin(x)}\]

OpenStudy (anonymous):

It should be LHS..

OpenStudy (anonymous):

Dividing numerator and denominator by 2 will start the solution..

OpenStudy (anonymous):

Yes that is the correct equation and the way I learned how to do it is that you have to make the left hand side equal the right. I will try dividing by 2.

OpenStudy (anonymous):

I still do not understand it.

OpenStudy (loser66):

ok, can we go from the RHS? it's easier

OpenStudy (loser66):

\[\sec (\frac{ \pi }{ 6 }- x ) = \frac{ 1 }{ \cos (\frac{ \pi }{ 6 } -x)}\]

OpenStudy (loser66):

= \[\frac{ 1 }{ \cos \frac{ \pi }{ 6 } \cos x + \sin \frac{ \pi }{ 6 } \sin x}\]

OpenStudy (loser66):

= \[\frac{ 1 }{ \sqrt{3} /2\cos x + \frac{ 1 }{ 2 }\sin x}\]

OpenStudy (loser66):

take 1/2 up, you got LHS

OpenStudy (anonymous):

Ok where did sin pi/6 come from? Wouldn't it be 1/(cos(pi/6)-cos x)?

OpenStudy (loser66):

nope, it's the formula cos (a- b) = cos a cos b+ sin a sin b your a is pi/6 your b is x

OpenStudy (loser66):

kapeech???

OpenStudy (anonymous):

So with the half in the denominator, do you multiply the numerator and denominator by 2?

OpenStudy (loser66):

you can do that. the same answer

OpenStudy (anonymous):

Oh ok thank you so much!! I get it now.

OpenStudy (loser66):

yw